OpenStudy (anonymous):

A roasted turkey is taken from an oven when its temperature has reached 185 Fahrenheit and is placed on a table in a room where the temperature is 75 Fahrenheit. (a) If the temperature of the turkey is 140 Fahrenheit after half an hour, what is its temperature after 45 minutes? Your answer is ? Fahrenheit. (b) When will the turkey cool to 100 Fahrenheit? Your answer is ? hours

OpenStudy (anonymous):

the temperature of the turkey goes down the longer it is on the table. In other words as t for time gets longer, temp gets less. This is indirect proportion. temp = k (1/time) k is a number we have to find. From part a. we know 140 = k(1/30) (30 mins = 1/2 an hour). NOTE we use time in minutes for all the rest of the equations here. so k= 140x30 = 4200 formula now reads temp=4200(1/time) =4200/time temp=4200/45 =93 +1/3 Fahrenheit. part b. formula was temp = 4200/time rearranging for time time=4200/temp =4200/100 =42 minutes.

OpenStudy (anonymous):

oh thanks i got it thought but this helps also!!:D